OpenStudy (anonymous):

I dont understand this can someone help will fan and medal! Cecilia owns two cars. The cost of maintenance and fuel for each car is a function of the distance the car covers (d). Cecilia’s total monthly expenses for the first car are given by the function f(d) = d^2 − 2d + 3, and her total monthly expenses for the second car are given by the function g(d) = d^2 + 5. The function that gives Cecilia’s total monthly expenses for both cars is h(d) =

OpenStudy (anonymous):

I will try my best to help you out. I apologize if I am wrong in advance. Okay, you have the expenses for one car and the expenses for the other car. Let's start with an analogy. You want to buy an apple and a sandwich. The apple costs $5, and the sandwich costs $11 . Apple=$5 Sandwich=$11 How do you find the TOTAL cost? You add them! :) Total expenses: $5 + $11= $16 So, your total expenses are 16 dollars. This process can be applied to this problem. f(d)= d^2 - 2d + 3 g(d)= d^2 + 5 In order to get the expenses for BOTH cars altogether, you will have to add the equations. h(d)=(d^2 - 2d + 3) + (d^2 + 5) Now, simplify by grouping like terms. h(d)=d^2 + d^2 - 2d +5+3 h(d)= 2d^2 - 2d +8 This should be your answer. :D

OpenStudy (anonymous):

Omg thank you so much this made me understand it so much better!! I really appreciate your help(: @ags2658

OpenStudy (anonymous):

Could you help me with three more?? @ags2658

OpenStudy (anonymous):

Sure! I can try my best on them. If you tell me what they are, I will get started. :D

OpenStudy (anonymous):

Thank you so much!! @ags2658